Suppose that you are the manager of a production department that uses 2400 boxes of rivets per year. The supplier quotes you a price of $9.50 per box for an order size of 199 boxes or less, a price of $9.00 per box for orders of 200 to 499 boxes, and a price of $8.00 per box for an order of 500 or more boxes. You assign a holding cost of h=20 percent of the price to this inventory. What order quantity would you use if the objective is to minimize total annual costs of holding, purchasing, and ordering? Assume ordering cost is $60/order. What is your new EOQ and the annual total cost? Use the table below. Show all calculations.

Qty. Range C-CostunitH-hC Range EOQ EOQ feasible? Candidate EOQ

What is the Total Cost Calculation for each Candidate EOQ?


What is the time between orders (months) for your order size?
Answer #1

EOQ = sqrt(2*Demand*ordering cost/Holding cost)

Demand = 2400

Ordering cost = 60

Qty range C H Range EOQ EOQ feasible Candidate EOQ
0-199 10 1.9 389.3 No
200-499 9 1.8 400.0 Yes 400
>500 8 1.6 424.3 No 500

Total cost = Demand/EOQ*Ordering cost + EOQ/2*H + C*Demand

Total cost for EOQ 400 is: 2400/400*60 + 400/2*1.8 + 9*2400 = 22,320

Total cost for EOQ 500 is: 2400/500*60 + 500/2*1.6 + 8*2400 = 19,888

So, 500 units is the EOQ due to lower total cost

Time between orders = 12/Demand*EOQ = 12/2400*500 = 2.5 months